Becky makes earrings and bracelets but finds out that she has not calculated the costs of the materials and finds that to know the costs and calculate an estimated budget she needs to figure out the following: For the earrings, you need 5 clear ornaments, 4 black ornaments, and ½ wire. For the bracelets, you need 10 clear ornaments and 3 black ornaments but you do not need wire. The costs of the earrings are 200 cents (per unit) and the costs of the bracelets are 300 cents (per unit) There are the following restrictions: 5 earrings and 10 bracelets must be greater than or equal to the minimum demand of 90 clients who are students. 4 earrings and 3 bracelets must be greater than or equal to the minimum demand of 48 clients who are teachers. 1 small earring (for effects of the problem would be ½ a normal earring ) must be greater than or equal to the minimum demand which are three girls (for effects of the problem 1 ½ of an adult).

Q: With the previous data: Minimize Z (cost in dollars).
Answer #1

Let E and B be the number of earrings and bracelets to be made

Minimize 200E + 300B (total cost in cents)

s.t.

5E + 10B >= 90

4E + 3B >= 48

0.5E + 0B >= 1.5

E, B >= 0

Solution by LINGO is following:

Optimal solution:

E = 8.4

B = 4.8

Total cost = 3120 cents = $ 31.20